# A Self-Improvable Polymer Discovery Framework Based on Conditional Generative Model

### v2

@@ 4.1 Conditional Generation
Generative models have gained substantial traction within the scientific community, showcasing their effectiveness and widespread adoption. However, within the realm of materials science, the primary challenge often lies not in simply generating physically reasonable candidates but rather in producing materials with specifically targeted properties, a process commonly referred to as inverse design. To illustrate this point, consider the case of polymer electrolytes where researchers aspire to discover novel polymers exhibiting enhanced ion conductivity. Historically, experimental development of new polymer electrolytes that can compete with PEO has been mainly focused on different trial errors, including altering the O/C ratio in the polymer backbone. However, this approach is inefficient, and the previous attempts were not completely successful.
In this study, we present an approach that leverages the transformer-based generative model, with a modified encoding approach to achieve conditional generation based on the desired property (see Experiment Setup for details). The encoding contains two layers of information: the SMILES representation of the repeating unit, and the class of the ion conductivity. When trained with more than 6000 data, the model not only learned how to generate the valid SMILES string as polymer repeating units but also the relationship between the repeating unit and the ion conductivity (please also see [8]). The outcome of the model is a generative batch of novel polymer repeating units, which exhibits a shifted distribution of ion conductivity with a notably higher mean value (0.89 mS/cm) when compared to the training set (0.06 mS/cm, excluding the added PEO polymers) (Figure 2). It is a remarkable improvement by a factor of 15.
@@ 4.2 Active learning
Presently, our method involves using Molecular Dynamics (MD) simulations to validate ionic conductivity in newly generated polymers. Polymers demonstrating higher ion conductivity are then added to our training dataset. We examined how the distribution of polymers evolved across various iterations of active learning. The immediate finding is that the average ion conductivity of polymers generated by a model after one retraining cycle exceeded that of the initial model by 13%, as shown in Figure 3. Further, the lowest ion conductivity reported from the 2nd iteration batch is 3.6 times that of the 1st iteration. Furthermore, in Figure 4, we highlight new repeating units discovered through this process, which show higher ion conductivity compared to those in the original training set. After the second iteration of this process, the number of new discoveries increased from 8 to 11. This progression indicates the model’s potential for ongoing enhancement, leading to increasingly effective outputs via a systematic feedback approach.
@@ 4.3 Discovered polymers
In Figure 4, we introduce 19 novel polymer repeating units whose ion conductivities, as confirmed by MD simulations, surpass that of PEO. It’s important to reiterate that PEO currently holds the record for the highest ion conductivity for dry polymers (around 1 mS/cm at 353K, and $Li^{+}.TFSI^{-}$ molality of 1.5 mol/kg) in this field, and despite extensive research, no superior polymer electrolytes have been found—until now.
Among these repeating units, multiple polyacetals stand out. Polyacetals are polymers with a high oxygen-to-carbon ratio, similar to PEO, which facilitates efficient lithium salt solvation and creates effective pathways for lithium ion transport. Notably, poly(1,3-dioxolane) (P(EO-MO))—a polyacetal with a repeating unit of 1,3-dioxolane included in our list—demonstrates considerable promise. Its MD-calculated ion conductivity is 1.515 (± 0.199) mS/cm. Although experimental measurements (referenced in [21]) show slightly lower conductivity for P(EO-MO) at 0.4 mS/cm, its potential as a polymer electrolyte candidate remains significant due to its improved ion transport efficacy.
Our findings also reveal that many candidate polymers incorporate new nitrogen and sulfur elements diverging from the traditional focus on polycarbonates which consist only of carbon and oxygen. Particularly, the polymer with the highest conductivity identified during our research has the repeating unit of ONCCOC. In molecular dynamics simulations, it demonstrated an average ionic conductivity approximately double that of polyethylene oxide (PEO). This evidence underscores the creativity of our models and significantly broadens the scope for future research in this area.
Figure 3: The MD simulation validation of polymers generated from two iterative candidate generations. The box plots show the mean and standard deviation in 5 MD simulations performed for each listed polymer.
@@
The calculated ionic conductivity, derived from the cluster Nernst-Einstein equation, arises from both ion diffusivity and clustering. To elucidate the mechanisms underpinning the superior ionic conductivity observed in generated polymers, we conducted a comparative analysis of conductivity, ion diffusivity, and concentrations of free ions between PEO and the polymers exhibiting enhanced performance, as illustrated in Fig. 5. In this context, "free ions" denote those not incorporated into any clusters and moving freely, with their concentration determined as an average across simulation durations. The analysis reveals that the augmented conductivity in the most effective polymers from the first two iterations is attributed to both an increase in ion diffusivity and a higher prevalence of free ion clusters. Interestingly, it was also noted that several of the developed polymers exhibit a more efficient dissociation of the $Li^{+}.TFSI^{-}$ salt compared to PEO, indicating a potential for improved ion transport properties.
Figure 5: Comparison of the ionic conductivity, ions’ diffusivity, and free ion concentration between PEO and the generated polymers with higher conductivity
@@ 5 Discussion and Conclusion
As mentioned earlier, the conditioned generative model, in its present state, is somewhat crude and unrefined. The model’s efficacy as the primary driver of polymer discovery underscores the need for further development and refinement. Data oversampling strategies, that have proven to be instrumental in the experiment, also need to be thoroughly and systematically investigated. Future efforts could focus on improving the current model or exploring alternative architectures, potentially enhancing the platform’s efficiency and accuracy in generating high-quality polymer candidates. Notably, developing generative algorithms for such specialized tasks presents its own set of challenges, particularly in evaluating their effectiveness. Unlike more straightforward metrics such as Mean Absolute Error (MAE) used in regression tasks, assessing the performance of generative models is less direct. A sister publication [8], which is released concurrently, delves deep into the benchmark of different generative model architectures in various generative tasks. It addresses the issue of evaluation by proposing a comprehensive and systematic methodology to benchmark different models for polymer SMILES code generation. This approach will significantly aid in comparing and optimizing generative models for polymer discovery.
@@
In this study, we selected ionic conductivity as the primary metric to identify new polymer electrolytes. However, as highlighted in previous studies [21, 22], a more holistic measure is efficacy, defined as the product of conductivity and cationic current fraction. This measure considers both conductivity and the mobility of the preferred charge (cation here) in the system. In our study, ion conductivity calculations were standardized at a salt concentration of 1.5 mol/kg, which is around the optimum for ion mobility for PEO. Moving forward, we plan to modify our evaluation criteria to include ion transport efficacy based on cation mobility and to establish a database encompassing various salt concentrations. Also, the synthesizability of the generated polymers needs to be considered as well, potentially by using it as an additional constraint for generation.
@@
Another vital aspect of our platform is the validation process. Currently, we employ classical molecular dynamics simulations, which are fast and effective but may lack robustness. The ultimate goal is to corroborate our findings through experimental validation. However, given the current limitations in high-throughput polymer synthesis and characterization, simulation-based validation remains the most feasible approach. Our future endeavors will likely involve validating candidates through simulations before proceeding with experimental trials. Furthermore, the reliance on classical force fields, which are typically tailored to specific polymer classes, raises concerns about their extrapolative power to other materials. There is a noticeable disagreement between the reported ionic conductivity for the poly(EO-MO) (OCCOC) from ref[21, 22] and our computed one, which could be a result of the artifact of the force field. To address this, one promising direction is the use of ab initio molecular dynamics (AIMD) simulations with methods like Density Functional Theory (DFT). While these simulations offer greater robustness, their computational expense is a significant barrier. A potential solution could be the development and use of machine-learned force fields that combine the accuracy of DFT with the efficiency of classical force fields. Additionally, experimental verification would be critical for the promising candidates.
@@
The workflow of our platform, which currently lacks full modularity and automation, also presents an opportunity for further improvement. Throughout the discovery campaign, the process has been devoid of human intervention, which lays the groundwork for developing a fully automated system. Future efforts will be directed toward enhancing the platform’s modularity, enabling compatibility with various generative models and validation methods. This improvement would not only streamline the discovery process but also expand the platform’s utility across different domains of polymer research.
@@ 6 Code and Data Availability
The dataset used to train the generative models can be accessed on the HTP-MD website: https://www.htpmd.matr.io/. The code for training the generative models can be found here: https://github.com/TRI-AMDD/PolyGen. The code for running the verification MD simulations can be found here: https://github.com/TRI-AMDD/htp_md

### v3

@@ 4.1 Conditional Generation
Generative models have gained substantial traction within the scientific community, showcasing their effectiveness and widespread adoption. However, within the realm of materials science, the primary challenge often lies not in simply generating physically reasonable candidates but rather in producing materials with specifically targeted properties, a process commonly referred to as inverse design. To illustrate this point, consider the case of polymer electrolytes, where researchers aspire to discover novel polymers exhibiting enhanced ion conductivity. Historically, experimental development of new polymer electrolytes that can compete with PEO has been mainly focused on different trial errors, including altering the O/C ratio in the polymer backbone [52, 53, 54]. However, this approach is inefficient, and the previous attempts were not completely successful in finding polymers that can replace the liquid electrolytes.
In this study, we present an approach that leverages the transformer-based generative model with a modified encoding approach to achieve conditional generation based on the desired property (see Experiment Setup for details). The encoding contains two layers of information: the SMILES representation of the repeating unit and the class of the ion conductivity. When trained with more than 6000 data, the model not only learned how to generate chemically valid SMILES string as polymer repeating units but also the relationship between the repeating unit and the ion conductivity (please also see [26]). The outcome of the model is a generative batch of polymer repeating units, which exhibits a shifted distribution of ion conductivity with a notably higher mean value (0.75 mS/cm) when compared to the training set (0.06 mS/cm, excluding the added PEO polymers) (Figure 2). It is a remarkable improvement by a factor of 12.
@@ 4.2 Active learning
Presently, our method involves using Molecular Dynamics (MD) simulations to estimate ionic conductivity in newly generated polymers. Polymers demonstrating higher ion conductivity are then added to our training dataset. We examined how the distribution of polymers evolved across various iterations of active learning. The immediate finding is that the average ion conductivity of polymers generated by a model after one retraining cycle exceeded that of the initial model by 15%, as shown in the box-and-whisker plot in Figure 3. Further, the lowest ion conductivity reported from the 2nd iteration batch is 3.6 times that of the 1st iteration. Furthermore, in Figure 4, we highlight new repeating units discovered through this process, which show higher ion conductivity compared to those in the original training set. This progression indicates the model’s potential for ongoing enhancement, leading to increasingly effective outputs via a systematic feedback approach. Other ion transport properties of the generated polymer in the two iterations, including density, ion diffusivity, ionic conductivity, and transference number, are listed in Supporting Information (Table.S5 and S6).
Although we observed an increase in the minimum and average ionic conductivity, the improvement in ionic conductivity and the number of polymers outperformed the baseline levels (7 polymers in each iteration), and there was, in fact, a decrease in the maximum ionic conductivity of generated polymers in the second iteration. We believe this decrease is due to the exhaustion of the limited search space. Since we are exploring linear chain homopolymers composed of only a few heavy atoms, the search space is extremely narrow. The limited search space is also the reason that only two iterations of exploration are performed. We believe that by extending the chemistry to a wider range of atom types and incorporating more diverse polymer structures (e.g., aromatic structures), a saturation of performance increase would occur at later stages. This highlights the necessity of developing more creative generative models that can extrapolate the search space, which has been the focus of other researchers’ studies [22].
@@ 4.3 Discovered polymers
In Figure 4, we introduce 14 generated polymer repeating units whose ionic conductivities, as confirmed by MD simulations, surpass that of PEO. It is important to reiterate that PEO currently holds the record for one of the highest ion conductivity in the form of dry (solid) polymers with around 1 mS/cm conductivity at 353 K, and $Li^{+}.TFSI^{-}$ molality of 1.5 mol/kg) and is still considered a baseline material in this field.
Among these polymer repeat units, several polyacetals stand out. Polyacetals are polymers with a high oxygen-to-carbon ratio, similar to PEO, which facilitates efficient lithium salt solvation and creates effective pathways for lithium ion transport. For example, poly(1,3-dioxolane) (P(EO-MO), *OCCOC*), a polyacetal with a repeating unit of 1,3-dioxolane, demonstrates considerable promise. Its MD-calculated ion conductivity is 1.515 (± 0.199) mS/cm. Although experimental measurements show lower conductivity for P(EO-MO) at 0.4 mS/cm, its potential as a polymer electrolyte candidate remains significant due to its improved ion transport efficacy [53]. Similarly, poly(diethylene oxide-alt-oxymethylene) (P(2EO-MO)), which appeared as *OCCOCCOC* in the first iteration, has been previously synthesized and investigated, showing slightly lower ionic conductivity of 1.1 mS/cm compared to PEO’s 1.5 mS/cm at 90°C [55]. Another notable polymer is polyethylene oxide-alt-trimethylene oxide (P(EO-TMO)), which appeared as *OCCOCCC* in the second iteration. Previous molecular dynamics simulations have supported the higher ionic conductivity of P(EO-TMO) compared to PEO [56], but we couldn’t find any experimentally measured ionic conductivity for this polymer.
Our research uncovered that many of the candidate polymers feature new elements like nitrogen and sulfur, marking a shift from the conventional focus on polycarbonates composed solely of carbon and oxygen. Notably, the polymer with the highest conductivity discovered in our study contains the repeating unit ONCCOC. Molecular dynamics simulations demonstrated that this polymer exhibits an average ionic conductivity roughly twice that of polyethylene oxide (PEO). Despite the absence of previous experimental or theoretical studies on this specific polymer, related research on polyethylenimine (PEI, CCN) polymer electrolytes [57, 58] and their blends with PEO [59, 60] has been documented. However, many of the generated polymers include bonds such as N-O or S-O and motifs like O-O-NH or O-NH-O, which present significant synthetic challenges due to their physically unrealistic structures. Nonetheless, the success of our framework in consistently generating polymers with improved ionic conductivity in MD simulations and the appearance of many of the generated candidates in the literature highlights the innovative potential of our models and greatly expands the possibilities for future research in this field.
Figure 3: Ionic conductivity of polymers generated from two iterative candidate generations computed from MD simulations. The box plots show the mean and standard deviation in 5 MD simulations performed for each listed polymer. The dashed and dotted lines show the mean and the standard deviation of ionic conductivity of PEO computed from MD simulation.
@@
The calculated ionic conductivity, derived from the cluster Nernst-Einstein equation, arises from both ion diffusivity and clustering. To elucidate the mechanisms underpinning the superior ionic conductivity observed in generated polymers, we conducted a comparative analysis of conductivity, ion diffusivity, and concentrations of free ions between PEO and the polymers exhibiting enhanced performance (see Fig.S1). In this context, "free ions" denote those not incorporated into any clusters and moving freely, with their concentration determined as an average across simulation durations. The analysis reveals that the augmented conductivity in the most effective polymers from the first two iterations is attributed to both an increase in ion diffusivity and a higher prevalence of free ion clusters. Interestingly, it was also noted that several of the developed polymers exhibit a more efficient dissociation of the $Li^{+}.TFSI^{-}$ salt compared to PEO, indicating a potential for improved ion transport properties.
Salt concentration is a crucial design parameter that influences ion pairing in polymer electrolytes. Both ionic conductivity and free ion concentration initially increase with higher salt concentration, but at very high concentrations, ion clusters of cations and anions form. These clusters reduce effective conductivity due to charge cancellation within each cluster [39]. The optimal salt concentration depends on how strongly polymer atoms coordinate with ions, which affects salt dissociation. Consequently, this optimal concentration varies for different polymers. Generally, a practical electrolyte system for lithium-ion batteries requires a moderate to high salt concentration to achieve enhanced ionic conductivity [61], mechanical strength [62], electrochemical stability [63], and solid electrolyte interphase (SEI) formation [64]. We examined the ionic conductivity of several top candidates across various salt concentrations at 353 K to illustrate this concept (see Fig. S2). The MD simulation results indicated that maximum conductivity occurs at slightly different salt molalities, generally around 1.5-2.0 mol/kg. Although the salt concentration in the training set used to generate the polymers in this study was 1.5 mol/kg, exploring this parameter further as a design factor is recommended for future research.
@@ 5 Discussion and Conclusion
As mentioned earlier, the conditioned generative model, in its present state, is somewhat crude and unrefined. The model’s efficacy as the primary driver of polymer discovery underscores the need for further development and refinement. Although the choice of using SMILES strings to represent materials allows us to apply the presented discovery framework to other cases, such as molecular discovery (e.g., liquid electrolytes), it also limits the information available to the model. SMILES strings are designed to represent molecules as linear sequences, which can be inadequate for capturing the complexity of branched or cross-linked polymer structures. However, it is possible to enhance SMILES strings with 3D structural information or to develop representations that are inherently aware of 3D structures. Data oversampling strategies that have proven to be instrumental in the experiment also need to be thoroughly and systematically investigated. Future efforts could focus on improving the current model or exploring alternative architectures, potentially enhancing the platform’s efficiency and accuracy in generating high-quality polymer candidates. Notably, developing generative algorithms for such specialized tasks presents its own set of challenges, particularly in evaluating their effectiveness. Unlike more straightforward metrics such as Mean Absolute Error (MAE) used in regression tasks, assessing the performance of generative models is less direct. A sister publication [26], which is released concurrently, delves deep into the benchmark of different generative model architectures in various generative tasks. It addresses the issue of evaluation by proposing a comprehensive and systematic methodology to benchmark different models for polymer SMILES code generation. This approach will significantly aid in comparing and optimizing generative models for polymer discovery.
@@
In this study, we selected ionic conductivity as the primary metric to identify new polymer electrolytes. However, as highlighted in previous studies [53, 52], a more holistic measure is efficacy, defined as the product of conductivity and cationic current fraction. This measure considers both conductivity and the mobility of the preferred charge (cation here) in the system. In our study, ion conductivity calculations were standardized at a salt concentration of 1.5 mol/kg, which is around the optimum for ion mobility for PEO. Moving forward, we plan to modify our evaluation criteria to include ion transport efficacy based on cation mobility and to establish a database encompassing various salt concentrations.
Additionally, to address the growing demands for multi-properties optimization for materials science problems, our platform can be adapted to evaluate crucial polymer properties such as glass transition temperature (Tg), bulk density, and mechanical properties. This would involve extending the current framework to incorporate these properties into the design criteria, allowing for a more comprehensive assessment of polymer candidates. Future research will explore incorporating these additional properties into the model input to enhance the platform’s capability in identifying polymers that not only exhibit superior ionic conductivity but also meet other critical performance metrics. We would also like to emphasize that beyond properties such as ionic conductivity, other critical factors are essential for successful polymer discovery, with synthesizability being one of the most important. Our platform can indeed incorporate considerations of synthesizability. For tools that quantify synthesizability, such as the SA score, we can generate and iterate on more synthesizable candidates in the same manner as we did for ion conductivity. For retrosynthesis planning tools [65], which propose synthesis routes for candidates, we can a) quantify synthesizability based on factors such as the number of plausible synthesis recipes, required synthesis conditions, and the kinetics of the routes, and use that as a condition; and/or b) enhance the feedback loop by incorporating real-world testing of the synthesis recipes.
@@
Another vital aspect of our platform is the computational evaluation process. Currently, we employ classical molecular dynamics simulations, which are fast and effective but may lack robustness. The ultimate goal is to corroborate our findings through experimental validation. However, given the current limitations in high-throughput polymer synthesis and characterization, simulation-based assessment remains the most feasible approach. Our future endeavors will likely involve evaluating candidates through simulations before proceeding with experimental trials. Furthermore, the reliance on classical force fields, which are typically tailored to specific polymer classes, raises concerns about their extrapolative power to other materials. There is a noticeable disagreement between the reported ionic conductivity for the poly(EO-MO) (OCCOC) from ref[53, 52] and our computed one, which could be a result of the artifact of the force field. To address this, one promising direction is the use of ab initio molecular dynamics (AIMD) simulations with methods like Density Functional Theory (DFT). While these simulations offer greater robustness, their computational expense is a significant barrier. A potential solution could be the development and use of machine-learned force fields that combine the accuracy of DFT with the efficiency of classical force fields. Additionally, experimental verification would be critical for the promising candidates.
@@
Although the workflow of our platform currently lacks full modularity and automation to provide more flexibility and expedite development, this approach presents an opportunity for further improvement. Throughout the discovery campaign, the process included minimal human intervention, laying the groundwork for developing a fully automated system. Future efforts will be directed toward enhancing the platform’s modularity, enabling compatibility with various generative models and evaluation methods. This improvement would not only streamline the discovery process but also expand the platform’s utility across different domains of polymer research.
@@ 6 Code and Data Availability
The dataset used to train the generative models can be accessed on the HTP-MD website: https://www.htpmd.matr.io/. The subset of HTP-MD dataset used for training has been included in the https://github.com/TRI-AMDD/PolyGen as htpmd-trainset.csv. The code for training the generative models can be found here: https://github.com/TRI-AMDD/PolyGen. The code for running the MD simulations can be found here: https://github.com/TRI-AMDD/PolyGen/tree/main/Example-simulation-files. Consistent with the trainset, the MD simulation trajectories were analyzed, and the ionic conductivity of the generated polymers has been computed using HTP-MD code available at https://github.com/TRI-AMDD/htp__\__md.